Here are book- and market-value balance sheets of the United Frypan Company (figures in $ millions):

Book-Value Balance Sheet
Net working capital $ 70 Debt $ 60
Long-term assets 30 Equity 40
$ 100 $ 100
Market-Value Balance Sheet
Net working capital $ 70 Debt $ 60
Long-term assets 150 Equity 160
$ 220 $ 220

Assume that MM’s theory holds except for taxes. There is no growth, and the $60 of debt is expected to be permanent. Assume a 21% corporate tax rate.

a. How much of the firm's market value is accounted for by the debt-generated tax shield? (Enter your answer in million rounded to 2 decimal places.)

b. What is United Frypan’s after-tax WACC if rDebt = 6.1% and rEquity = 16.9%? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)

c. Now suppose that Congress passes a law that eliminates the deductibility of interest for tax purposes after a grace period of 5 years. What will be the new value of the firm, other things equal? Assume a borrowing rate of 6.1%. (Do not round intermediate calculations. Enter your answer in million rounded to 2 decimal places.)
Answer #1

Firm value always refer to market value ,

Answer a) Market value of tax shield = tax rate * Debt value = 21%*60= $12,6 Million

Answer b) WACC = Wd *Rd*(1-T)+We*Re

WACC=[60/220*0.061*(1-0.33)+160/220*0.169]=0.13406 i.e. 13.41%

Answer c) Annual Tax shield = Tax * Interest Expenses = 21%*6.1%*60= 0.7686 million.

PV of Tax shield in five years = 0.7686* Annuity factor(6.1%,5)= 3.229

The total value of firm fall by = 12.6 -3.229 = $ 9.371 million

New value of firm = 220 -9.371= $ 210.629 million i.e. $ 210.63 million.